OpenStudy (anonymous):

Convert the following complex number into its polar representation:

OpenStudy (anonymous):

\[2-2\sqrt{3}i\]

OpenStudy (anonymous):

Please help!! Live online test!!!!

OpenStudy (anonymous):

@zzr0ck3r can you help me?

OpenStudy (anonymous):

@wio ????

OpenStudy (anonymous):



OpenStudy (anonymous):

Answer is \[ r(\cos(\theta) +i\sin(\theta)) \]

OpenStudy (anonymous):

Just need to find \(r\) and \(\theta\).

OpenStudy (anonymous):

ok thanks a bunch, but how do you do that?

OpenStudy (anonymous):

Okay, first of all, it's a right triangle. Do you really not know how to solve for \(r\)?!

OpenStudy (anonymous):

the answer is A right!

OpenStudy (anonymous):

R=square root of A^2+B^2

OpenStudy (anonymous):

Is it A @wio

OpenStudy (anonymous):

Yes...

OpenStudy (anonymous):

Thanks!!

OpenStudy (anonymous):

@wio i have another question which is similar